3. Phoenix produces dividends in four consecutive years of 0.32, 0.62, 0.82, and 0.92 respectively. The dividend in year five is estimated to be 1.05 and should grow in perpetuity at 3% thereafter. Given a discount rate of 10%, what is the price of the stock today?
Answer #1

Value after year 5=(D5*Growth rate)/(Discount rate-Growth rate)

=(1.05*1.03)/(0.1-0.03)=$15.45

Hence current price of the stock=Future dividends*Present value of discounting factor(10%,time period)

=0.32/1.1+0.62/1.1^2+0.82/1.1^3+0.92/1.1^4+1.05/1.1^5+15.45/1.1^5

which is equal to

$12.29(Approx)